# Motion-R1: Chain-of-Thought Reasoning and Reinforcement Learning for Human Motion Generation

### v3

@@ 3.4 Training Strategy
3.4.1 Cold-Start Training with CoT
Inspired by the success of DeepSeek-R1-Zero deepSeekR1IncentivizingReasoning2025 in solving mathematical reasoning tasks via pure RL, we initially explore end-to-end RL training for our model, aiming to jointly induce CoT reasoning and motion token generation solely from reward signals. However, this approach proves highly unstable in our setting: the model frequently fails to generate coherent CoT sequences and, more critically, produces motion tokens that are either semantically misaligned or syntactically invalid.
These challenges may be partially explained by two fundamental differences between our task and the conventional domains of symbolic reasoning. First, unlike the short-form outputs (e.g., numerical answers) in math problems, motion generation requires long, structured sequences of discrete tokens. Second, these motion tokens are newly introduced symbols in the model’s vocabulary, whose embeddings are insufficiently trained to bridge the modality gap between natural language and motion representations. As a result, the model struggles to interpret and generate them correctly.
To address these issues, we adopt a cold-start strategy based on supervised fine-tuning. Leveraging the curated dataset, comprising aligned triplets of task descriptions, CoT reasoning process, and motion sequences, we fine-tune the language model to bootstrap its ability to generate both structured CoT explanations and valid motion outputs. This supervised initialization constrains the model to learn the expected format and semantics of both reasoning and motion, providing a stable and effective foundation for subsequent policy optimization via RL.
3.4.2 GRPO-based Reinforcement Learning
To bridge the gap between natural language semantics and motion synthesis, we formulate text-to-motion generation as an RL problem and adopt GRPO shao2024deepseekmathpushinglimitsmathematical to optimize motion generation policies without relying on explicit reward functions or human-labeled preferences.
Unlike traditional RL approaches such as PPO schulman2017proximal or DPO rafailov2023direct , which require value networks or pairwise preference data, GRPO is a lightweight and scalable alternative. It leverages intra-group ranking feedback to optimize the generation policy by encouraging better samples within each group. This eliminates the need for training an additional critic model and reduces the overall computational overhead.
@@ 3.4.2 GRPO-based Reinforcement Learning
where $\varepsilon$ and $\beta$ are hyperparameters controlling the clipping range and KL regularization strength, respectively. $\pi_{\text{ref}}$ denotes the reference policy model. The normalized advantage term is given by $\hat{A}_{i}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$. This formulation enables stable and efficient policy optimization, incorporating semantic- and motion-aware reward shaping tailored specifically for the text-to-motion generation task.
To better guide policy learning and improve alignment between textual prompts and generated motions, we introduce a multi-reward mechanism.
Format Reward. To ensure that the content generated by the model has a resolvable structure, we introduce Format Reward $r_{\text{format}}$. This reward detects through regularization expressions whether the generated results strictly follow the predefined format: <think>{CoT}</think><Motion>{Motion tokens}</Motion>. Here, curly braces and their enclosed content denote placeholders, representing the generated CoT and the corresponding motion tokens, respectively. If the generated content exactly matches this required format, we assign a reward of 1. Otherwise, the reward is set to 0.
Motion Similarity Reward. To encourage temporal and spatial coherence between the generated motion $\hat{\mathbf{m}}$ and the ground-truth motion $\mathbf{m}$, we adopt a motion similarity reward $r_{\text{motion}}$. Specifically, we employ a pre-trained motion encoder $f_{\text{motion}}$, as introduced in Guo_2022_CVPR , to obtain feature representations of both motions. The reward is computed as the cosine similarity between their embeddings:
@@
Semantic Similarity Reward. To ensure that the generated motion semantically corresponds to the input textual description $T$, we introduce the semantic similarity reward $r_{\text{semantic}}$. This reward measures the alignment between the motion embedding and the language embedding in a shared latent space, both extracted using pre-trained encoders $f_{\text{motion}}$ and $f_{\text{text}}$, respectively (also from Guo_2022_CVPR ). The reward is defined as:
@@
These reward signals are not directly used in scalar form but are reflected through groupwise preference ranking in GRPO.
@@ 4.1 Experimental Settings
Datasets We evaluate our method on two widely-used benchmarks: HumanML3D Guo_2022_CVPR and KIT-ML plappertKITMotionLanguageDataset2016 . HumanML3D contains $14,616$ motions and $44,970$ textual descriptions, sourced AMASS AMASS2019 and HumanAct12 guo2020action2motion . KIT-ML includes $3,911$ motion clips and $6,278$ descriptions, derived from KIT kit2015 and CMU CarnegieMellonUniversity datasets. Both datasets follow standard train/val/test splits.
Table 1: Quantitative results of Motion-R1 on HumanML3D. The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.
@@
Top 1 Top 2 Top 3
MDM (tevet2022human, ) $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $\textbf{2.799}^{\pm.072}$
MLD (chenExecutingYourCommands2023, ) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
MotionDiffuse (zhang2023motiondiffuse, ) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
T2M (Guo_2022_CVPR, ) $0.457^{\pm.002}$ $0.559^{\pm.007}$ $0.740^{\pm.003}$ $1.067^{\pm.002}$ $3.340^{\pm.008}$ $9.188^{\pm.002}$ $2.090^{\pm.083}$
TM2T (guoTM2TStochasticTokenized2022, ) $0.424^{\pm.003}$ $0.618^{\pm.003}$ $0.729^{\pm.002}$ $1.501^{\pm.017}$ $3.467^{\pm.011}$ $8.589^{\pm.076}$ $\underline{2.424}^{\pm.093}$
T2M-GPT (zhangT2MGPTGeneratingHuman2023, ) $0.491^{\pm.003}$ $0.680^{\pm.003}$ $0.775^{\pm.002}$ $\underline{0.116}^{\pm.004}$ $3.118^{\pm.011}$ $9.761^{\pm.081}$ $1.856^{\pm.011}$
MotionGPT (jiangMotionGPTHumanMotion2023, ) $0.492^{\pm.003}$ $0.681^{\pm.003}$ $0.778^{\pm.002}$ $0.232^{\pm.008}$ $3.096^{\pm.008}$ $9.528^{\pm.071}$ $2.008^{\pm.084}$
MoMask guo2023momask $\textbf{0.521}^{\pm.002}$ $\underline{0.713}^{\pm.002}$ $\underline{0.807}^{\pm.002}$ $\textbf{0.045}^{\pm.002}$ $\underline{2.958}^{\pm.008}$ ${9.620}^{\pm.064}$ $1.241^{\pm.040}$
MotionChain (jiang2023motionchain, ) $0.504^{\pm.003}$ $0.617^{\pm.002}$ $0.790^{\pm.003}$ $0.248^{\pm.009}$ $3.033^{\pm.010}$ $9.470^{\pm.075}$ $1.727^{\pm.014}$
MotionLLM wang2024motionagent $0.515^{\pm.004}$ $0.691^{\pm.003}$ $0.801^{\pm.004}$ $0.230^{\pm.009}$ $2.967^{\pm.020}$ $9.908^{\pm.102}$ $2.142^{\pm.014}$
MotionGPT-2 wangMotionGPT2GeneralPurposeMotionLanguage2024 $0.496^{\pm.002}$ $0.691^{\pm.003}$ $0.782^{\pm.004}$ $0.191^{\pm.004}$ $3.080^{\pm.013}$ $9.860^{\pm.026}$ $2.137^{\pm.022}$
@@
Evaluation Metrics Following standard protocols Guo_2022_CVPR ; zhangT2MGPTGeneratingHuman2023 ; wang2024motionagent , we report R-Precision@1/2/3, FID, Diversity, MM-Dist, and MModality. These metrics respectively evaluate retrieval accuracy, distributional realism, sample diversity, text-motion alignment, and one-to-many generation ability. Note that more details about the evaluation metrics can be found in the supplementary material.
4.2 Implementation Details
For constructing the MotionCoT Data Engine, we utilize DeepSeek-R1 deepSeekR1IncentivizingReasoning2025 (API Version: 250120) to generate detailed CoT explanations corresponding to complex motion descriptions. These structured reasoning traces serve as intermediate semantic representations that bridge textual instructions and action planning. In Motion-R1, we adopt Qwen-2.5-3B-Instruct qwen25 as the backbone language model for generating CoT sequences and guiding motion synthesis. Qwen-2.5 is selected for its strong performance in multi-step reasoning and compact model size, which ensures efficient training within our RL pipeline. The generated CoTs are subsequently used as inputs to the motion generator and as targets for reward-based optimization via GRPO. All of our experiments are conducted on NVIDIA H20 GPUs.
@@ 4.3 Quantitative Evaluation
We evaluate the performance of Motion-R1 on HumanML3D and KIT-ML datasets, comparing it with state-of-the-art methods, ranging from VAE-based to diffusion-based models. For the diffusion-based model, we select MDM tevet2022human , MLD chenExecutingYourCommands2023 , MotionDiffuse zhang2023motiondiffuse . For the VAE-based model, we choose T2M Guo_2022_CVPR , TM2T guoTM2TStochasticTokenized2022 , T2M-GPT zhangT2MGPTGeneratingHuman2023 , MotionGPT jiangMotionGPTHumanMotion2023 , MoMask guo2023momask , MotionChain jiang2023motionchain , MotionLLM wang2024motionagent and MotionGPT-2 wangMotionGPT2GeneralPurposeMotionLanguage2024 .
We utilize the same evaluation settings as in previous works Guo_2022_CVPR ; wang2024motionagent , ensuring a fair comparison. For each method, we report the average results across 20 times (except MModality 5 times) to obtain a $95\%$ confidence interval. The quantitative results on HumanML3D are presented in Table 1. As shown in Table 1, Motion-R1 achieves state-of-the-art performance across multiple evaluation metrics on the HumanML3D dataset. In terms of R-Precision, Motion-R1 attains an R-Precision@1 of $0.515$, which is on par with the best-performing method (MoMask guo2023momask ), while achieving the highest R-Precision@2 ($0.719$) and R-Precision@3 ($0.818$), demonstrating superior semantic alignment and reasoning capabilities in motion-to-text retrieval. In terms of motion quality, Motion-R1 achieves a low FID of $0.201$, which is competitive with recent strong baselines such as MotionGPT-2 ($0.191$) and MotionGPT ($0.232$), indicating high realism and fidelity of the generated motions. Our method outperforms all baselines in MM-Dist ($2.854$), highlighting better alignment in the shared embedding space of motion and language. Moreover, Motion-R1 significantly outperforms all competing methods on the Diversity metric, reaching a score of $10.026$, suggesting that the model effectively captures the one-to-many nature of text-to-motion mappings. On the MModality metric, our approach achieves a strong score of $2.317$, which is competitive with other top methods, further validating the richness of generated motions conditioned on the same textual input.
Figure 4: Visualization comparisons with MotionLLM wang2024motionagent on in-distribution and out-of-distribution prompts.
@@ 4.4 Qualitative Analysis and Visualization
To complement the quantitative evaluation, we provide qualitative comparisons between Motion-R1 and existing state-of-the-art methods. We divide this section into two parts: (1) in-distribution prompts from standard benchmarks, and (2) out-of-distribution instructions that require compositional reasoning and generalization. These visualizations highlight the advantages of Motion-R1 in producing semantically coherent, diverse, and controllable motion sequences.
In-Distribution Visualization. We first visualize motion generation results on descriptions drawn from the HumanML3D and KIT-ML test sets. These examples cover various action types, including short actions and multi-step descriptions. We compare Motion-R1 against representative baselines such as MoMask and MotionGPT. As shown in Figure 4 (left), Motion-R1 produces smooth, well-structured sequences for simple and multi-step instructions. For example, given “A person walks in a circle.” our method generates a continuous loop with clear orientation change and natural timing, while MotionLLM wang2024motionagent fails to complete a full circle or exhibits abrupt terminations. This illustrates Motion-R1’s better alignment with spatial and temporal semantics, even in standard settings.
Table 2: Ablation study on HumanML3D. CoT, $R_{\text{sem}}$, and $R_{\text{motion}}$ denote the Chain-of-Thought decomposition, the semantic similarity reward, and the motion similarity reward, respectively. The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.

### v4

@@ 3.4 Training Strategy
3.4.1 Cold-Start Training with Decomposed CoT
Inspired by DeepSeek-R1-Zero Guo et al. (2025), we first attempt end-to-end RL training to induce decomposed CoT reasoning and motion generation purely from reward signals. Yet this setting proves unstable, as the model often fails to produce coherent reasoning or valid motion tokens. This instability stems from two factors: motion generation requires long, structured sequences rather than short symbolic outputs, and motion tokens are newly introduced symbols with insufficiently trained embeddings to bridge the modality gap.
To mitigate these issues, we adopt a cold-start strategy based on supervised fine-tuning. Leveraging curated triplets of captions, decomposed CoT traces, and motions, we bootstrap the model’s capacity to produce structured reasoning and valid motion outputs, providing a stable foundation for subsequent RL optimization.
3.4.2 GRPO-based Training with RL Binding
To enhance alignment between textual instructions and generated motions without incurring costly human annotations, RL Binding formulates text-to-motion generation as a reinforcement learning problem, employing GRPO Shao et al. (2024) to optimize the generation policy efficiently.
@@ 3.4.2 GRPO-based Training with RL Binding
where $\varepsilon$ and $\beta$ are hyperparameters controlling the clipping range and KL regularization strength, respectively. $\pi_{\text{ref}}$ denotes the reference policy model. The normalized advantage term is given by $\hat{A}_{i}=\frac{r_{i}-\text{mean}(\mathbf{r})}{\text{std}(\mathbf{r})}$, providing a normalized signal that facilitates stable and efficient policy updates.
The reward mechanism in RL Binding focuses on motion similarity and semantic alignment, ensuring that generated motions are temporally coherent and semantically faithful to the textual input. A format reward is also included to maintain structural validity, collectively enabling effective reinforcement learning without costly human annotations.
Format Reward. To ensure that the content generated by the model has a resolvable structure, we introduce Format Reward $r_{\text{format}}$. This reward detects through regularization expressions whether the generated results strictly follow the predefined format:
<think>{Decomposed CoT}</think><Motion>{Motion tokens}</Motion>.
Here, curly braces and their enclosed content denote placeholders, representing the generated CoT and the corresponding motion tokens, respectively. If the generated content exactly matches this required format, we assign a reward of 1. Otherwise, the reward is set to 0.
Motion Similarity Reward. To enforce temporal and spatial consistency, RL Binding computes a motion similarity reward $r_{\text{motion}}$ between the generated motion $\hat{\mathbf{m}}$ and the ground-truth motion $\mathbf{m}$. A pre-trained motion encoder $f_{\text{motion}}$ Guo et al. (2022a) is used to extract feature embeddings, and the reward is defined as the cosine similarity between these embeddings:
@@
Semantic Similarity Reward. To ensure that the generated motion semantically corresponds to the input textual description $T$, RL binding evaluates the semantic similarity reward $r_{\text{semantic}}$, which measures the alignment between the motion embedding and the language embedding in a shared latent space. Both embeddings are extracted using pre-trained encoders $f_{\text{motion}}$ and $f_{\text{text}}$, respectively (also from Guo et al. (2022a)). The reward is defined as:
@@
These rewards are integrated into GRPO via groupwise preference ranking, enabling RL Binding to simultaneously enforce motion realism and semantic fidelity.
@@ 4.1 Experimental Settings
Datasets We evaluate our method on two widely-used benchmarks: HumanML3D Guo et al. (2022a), KIT-ML Plappert et al. (2016) and BABEL Punnakkal et al. (2021). HumanML3D contains $14,616$ motions and $44,970$ textual descriptions, sourced AMASS Mahmood et al. (2019) and HumanAct12 Guo et al. (2020). KIT-ML includes $3,911$ motion clips and $6,278$ descriptions, derived from KIT Mandery et al. (2015) and CMU CMU Graphics Lab datasets. BABEL provides over 43 hours of motion capture sequences with 28k sequence-level and 63k frame-level action annotations from AMASS Mahmood et al. (2019).
Evaluation Metrics Following standard protocols Guo et al. (2022a); Zhang et al. (2023a); Wu et al. (2024), we report R-Precision@1/2/3, FID, Diversity, MM-Dist, and MModality. These metrics respectively evaluate retrieval accuracy, distributional realism, sample diversity, text-motion alignment, and one-to-many generation ability.
Table 1: Quantitative results of Motion-R1 on HumanML3D Guo et al. (2022a) and KIT-ML Plappert et al. (2016). The evaluations are conducted 20 times to obtain a $95\%$ confidence interval. Best results are highlighted in bold and the second best in underline.
@@
Top 1
Top 2
Top 3
HumanML3D
MDM (Tevet et al., 2023) $0.320^{\pm.005}$ $0.498^{\pm.004}$ $0.611^{\pm.007}$ $0.544^{\pm.044}$ $5.566^{\pm.027}$ $9.559^{\pm.086}$ $\textbf{2.799}^{\pm.072}$
MLD (Chen et al., 2023) $0.481^{\pm.003}$ $0.673^{\pm.003}$ $0.772^{\pm.002}$ $0.473^{\pm.013}$ $3.196^{\pm.010}$ $9.724^{\pm.082}$ $2.413^{\pm.079}$
MotionDiffuse (Zhang et al., 2022a) $0.491^{\pm.001}$ $0.681^{\pm.001}$ $0.782^{\pm.001}$ $0.630^{\pm.001}$ $3.113^{\pm.001}$ $9.410^{\pm.049}$ $1.553^{\pm.042}$
T2M (Guo et al., 2022a) $0.457^{\pm.002}$ $0.559^{\pm.007}$ $0.740^{\pm.003}$ $1.067^{\pm.002}$ $3.340^{\pm.008}$ $9.188^{\pm.002}$ $2.090^{\pm.083}$
TM2T (Guo et al., 2022b) $0.424^{\pm.003}$ $0.618^{\pm.003}$ $0.729^{\pm.002}$ $1.501^{\pm.017}$ $3.467^{\pm.011}$ $8.589^{\pm.076}$ $\underline{2.424}^{\pm.093}$
T2M-GPT (Zhang et al., 2023a) $0.491^{\pm.003}$ $0.680^{\pm.003}$ $0.775^{\pm.002}$ $\underline{0.116}^{\pm.004}$ $3.118^{\pm.011}$ $9.761^{\pm.081}$ $1.856^{\pm.011}$
MotionGPT (Jiang et al., 2023) $0.492^{\pm.003}$ $0.681^{\pm.003}$ $0.778^{\pm.002}$ $0.232^{\pm.008}$ $3.096^{\pm.008}$ $9.528^{\pm.071}$ $2.008^{\pm.084}$
MoMask Guo et al. (2024) $\textbf{0.521}^{\pm.002}$ $\underline{0.713}^{\pm.002}$ $\underline{0.807}^{\pm.002}$ $\textbf{0.045}^{\pm.002}$ $\underline{2.958}^{\pm.008}$ ${9.620}^{\pm.064}$ $1.241^{\pm.040}$
MotionChain (Jiang et al., 2024) $0.504^{\pm.003}$ $0.617^{\pm.002}$ $0.790^{\pm.003}$ $0.248^{\pm.009}$ $3.033^{\pm.010}$ $9.470^{\pm.075}$ $1.727^{\pm.014}$
MotionLLM Wu et al. (2024) $0.515^{\pm.004}$ $0.691^{\pm.003}$ $0.801^{\pm.004}$ $0.230^{\pm.009}$ $2.967^{\pm.020}$ $\underline{9.908}^{\pm.102}$ $2.142^{\pm.014}$
MotionGPT-2 Wang et al. (2024) $0.496^{\pm.002}$ $0.691^{\pm.003}$ $0.782^{\pm.004}$ $0.191^{\pm.004}$ $3.080^{\pm.013}$ $9.860^{\pm.026}$ $2.137^{\pm.022}$
@@
KIT-ML
TM2T (Guo et al., 2022b) $0.280^{\pm.005}$ $0.463^{\pm.006}$ $0.587^{\pm.005}$ $3.599^{\pm.153}$ $4.591^{\pm.026}$ $9.473^{\pm.117}$ $\textbf{3.292}^{\pm.081}$
T2M (Guo et al., 2022a) $0.361^{\pm.006}$ $0.559^{\pm.007}$ $0.681^{\pm.007}$ $3.022^{\pm.107}$ $3.488^{\pm.028}$ $10.720^{\pm.143}$ $2.052^{\pm.107}$
MDM (Tevet et al., 2023) $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $\underline{0.497}^{\pm.021}$ $9.191^{\pm.022}$ $10.850^{\pm.109}$ $1.907^{\pm.214}$
MotionDiffuse (Zhang et al., 2022a) $\underline{0.417}^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $\underline{2.958}^{\pm.005}$ $\underline{11.100}^{\pm.143}$ $0.730^{\pm.013}$
MLD (Chen et al., 2023) $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $0.404^{\pm.027}$ $3.204^{\pm.027}$ $10.800^{\pm.117}$ $2.192^{\pm.071}$
T2M-GPT (Zhang et al., 2023a) $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.023}$ $10.920^{\pm.108}$ $1.570^{\pm.039}$
AttT2M Zhong et al. (2023) $0.413^{\pm.005}$ $\underline{0.632}^{\pm.006}$ $\underline{0.751}^{\pm.006}$ $0.870^{\pm.039}$ $3.039^{\pm.021}$ $10.960^{\pm.123}$ $\underline{2.281}^{\pm.047}$
MotionLLM Wu et al. (2024) $0.409^{\pm.006}$ $0.624^{\pm.007}$ $0.750^{\pm.005}$ $0.781^{\pm.026}$ $\textbf{2.982}^{\pm.022}$ $\textbf{11.407}^{\pm.103}$
Motion-R1 $\textbf{0.431}^{\pm.003}$ $\textbf{0.638}^{\pm.002}$ $\textbf{0.761}^{\pm.003}$ $\textbf{0.287}^{\pm.004}$ $3.196^{\pm.040}$ $10.875^{\pm.052}$ $2.262^{\pm.014}$
Implementation Details We utilize DeepSeek-R1 Guo et al. (2025) to generate structured reasoning traces for complex motion descriptions, serving as intermediate representations that bridge textual prompts and motion tokens. Within Motion-R1, we adopt Qwen-2.5-3B-Instruct Team (2024) as the backbone model for its strong multi-step reasoning and efficient size, facilitating stable RL training. The generated CoTs condition motion synthesis and provide targets for GRPO optimization. Experiments are conducted on NVIDIA H20 GPUs. Additional experimental details are provided in Appendix A.4.
@@ 4.2 Quantitative Evaluation
We evaluate the performance of Motion-R1 on HumanML3D and KIT-ML datasets, comparing it with state-of-the-art methods, ranging from VAE-based to diffusion-based models. For the diffusion-based model, we select MDM Tevet et al. (2023), MLD Chen et al. (2023), MotionDiffuse Zhang et al. (2022a). For the VAE-based model, we choose T2M Guo et al. (2022a), TM2T Guo et al. (2022b), T2M-GPT Zhang et al. (2023a), MotionGPT Jiang et al. (2023), MoMask Guo et al. (2024), MotionChain Jiang et al. (2024), MotionLLM Wu et al. (2024) and MotionGPT-2 Wang et al. (2024).
We follow the same evaluation settings as prior work Guo et al. (2022a); Wu et al. (2024) and the BABEL setup from Zhuo et al. (2024). Quantitative results on HumanML3D and KIT-ML are shown in Table 1. Results on BABEL are provided in the Appendix A.5.
On HumanML3D, Motion-R1 attains strong and balanced performance across semantic and motion-quality metrics. It achieves R-Precision@1/2/3 = 0.515 / 0.719 / 0.818, with the latter two values being the best in the table and R-Precision@1 at near-top level. In terms of realism, Motion-R1 yields a competitive FID of 0.201, comparable to recent strong baselines (e.g., MotionGPT-2: 0.191, MotionGPT: 0.232). Motion-R1 also attains the lowest MM-Dist (2.854), indicating superior alignment in the joint motion–language embedding space, and the highest Diversity score (10.026), suggesting better modeling of the one-to-many nature of text-to-motion. Its MModality score (2.317) is likewise competitive with top methods. Together, these results indicate that the combination of decomposed CoT supervision and RL Binding yields motions that are both semantically aligned and high-fidelity.
On KIT-ML, Motion-R1 consistently leads the comparators on retrieval and fidelity metrics: R-Precision@1/2/3 = 0.431 / 0.638 / 0.761 (all best), and FID = 0.287 (best). While MM-Dist (3.196) is not the lowest in the table, Motion-R1 maintains strong Diversity (10.875) and solid MModality (2.262), demonstrating robust generalization across a dataset with different statistics. Overall, Motion-R1 provides a favorable trade-off between semantic alignment, motion realism, and diversity on both benchmarks, empirically supporting the effectiveness of the Decomposed CoT Data Engine and RL Binding.
@@ 4.3 Qualitative Analysis and Visualization
To complement the quantitative evaluation, we provide qualitative comparisons between Motion-R1 and existing state-of-the-art methods. We divide this section into two parts: (1) in-distribution prompts from standard benchmarks, and (2) out-of-distribution instructions that require compositional reasoning and generalization. These visualizations highlight the advantages of Motion-R1 in producing semantically coherent, diverse, and controllable motion sequences. Note that more visualizations are provided in Appendix A.2.
In-Distribution Visualization. We visualize results and compare Motion-R1 with MotionLLM as shown in Figure 3 (left), Motion-R1 generates smooth, coherent sequences that respect spatial and temporal semantics. For instance, for “A person walks in a circle,” it produces a continuous loop with natural timing, whereas MotionLLM Wu et al. (2024) often fails to complete the circle or shows abrupt stops.
Out-of-Distribution Visualization. To evaluate generalization beyond the training distribution, we present qualitative comparisons under two out-of-distribution captions (Figure 3, middle, right). The first prompt describes a multi-stage reaction: “After hearing a loud noise, a person turned around, stepped back cautiously with hands raised defensively, and then slowly approached.” Motion-R1 generates a temporally coherent sequence, clearly separating reaction, retreat, and re-approach, whereas MotionLLM merges steps or omits defensive gestures. The second prompt, “A person is serving in badminton,” involves abstract semantics and task-specific dynamics. Motion-R1 produces a plausible serve-like motion with arm lift, body shift, and forward striking gesture, while MotionLLM generates repetitive or generic movements.
These examples show that Motion-R1 generalizes well to unseen instructions through explicit CoT-style reasoning, whereas baseline methods struggle with long-horizon dependencies and abstract actions.
4.4 Ablation Study
To assess the contribution of each component in our framework, we conduct an ablation study on the HumanML3D dataset, as shown in Table 2. Specifically, we evaluate the impact of three key components: Decomposed CoT Data Engine and RL Binding (the semantic similarity reward ($R_{\text{sem}}$), and the motion similarity reward ($R_{\text{motion}}$).
All ablated variants achieve comparable performance, showing the robustness of our design. Using the Decomposed CoT Data Engine alone yields the weakest results (R-Precision Top-1: $0.340{\pm}.002$, FID: $0.530{\pm}.015$), while adding either $R_{\text{sem}}$ or $R_{\text{motion}}$ significantly improves alignment (Top-1: $0.483{\pm}.002$, FID: $0.281{\pm}.008$).
With all components combined, the model reaches the best or second-best performance across metrics, including the lowest FID ($\mathbf{0.201}{\pm}.004$) and highest R-Precision (Top-1: $\mathbf{0.515}{\pm}.003$). These results confirm the complementary roles of structured reasoning and semantic/motion-level rewards in producing diverse, semantically faithful, and high-quality motion.
Table 2: Ablation study on HumanML3D Guo et al. (2022a). CoT, $R_{\text{sem}}$, and $R_{\text{motion}}$ denote Decomposed CoT Data Engine, the semantic similarity reward, and the motion similarity reward, respectively. Best results are highlighted in bold and the second best in underline.